# RoboTwin 2.0: A Scalable Data Generator and Benchmark with Strong Domain Randomization for Robust Bimanual Robotic Manipulation

### v1

@@ 4.2 Evaluating Efficiency with and without Adaptive Grasping
To evaluate the effectiveness of our embodiment-aware grasp augmentation strategy, we measure the task success rate of automated data collection across 50 RoboTwin 2.0 tasks on five different robot embodiments. As shown in Table 4.2, we compare our RoboTwin 2.0 pipeline against the RoboTwin 1.0 baseline, which lacks diverse grasping and candidate augmentation. Results show that our method improves success rates, particularly for robots with constrained planning spaces, achieving an average improvement of 8.3% across all embodiments. Specifically, for high-DoF arms with large reachable workspaces, such as Franka and UR5 (7-DoF), success rates remain largely unchanged, indicating limited benefit when the robot already has sufficient kinematic flexibility. However, for lower-DoF platforms such as Aloha-AgileX, Piper, and ARX-X5 (6-DoF), our method leads to substantial gains of 13.5%, 22.7%, and 5.6%, respectively. These results demonstrate that our approach provides additional feasible grasp options that effectively mitigate the planning limitations of low-DoF manipulators. Success rates for all tasks can be found in Appendix H.
@@
To assess the impact of RoboTwin 2.0 data on improving model robustness to environmental perturbations, we design an ablation study based on the VLA framework. We independently collect a total of 9,600 expert trajectories across 32 tasks (300 per task) under two distinct settings: RoboTwin 2.0 with domain randomization and RoboTwin 1.0, which is RoboTwin 2.0 without randomization. Each dataset is used separately to fine-tune pretrained models: RDT for 10,000 steps and Pi0 for 30,000 steps, using offline imitation learning. To evaluate generalization, we select five unseen tasks and collect 50 new expert demonstrations under clean (non-randomized) conditions for single-task training and fine-tuning. As a control group, we include non-pretrained versions of ACT, DP, RDT, and Pi0. This setup enables a controlled evaluation of the effects of domain-randomized data on policy robustness in previously unseen environments. All experimental configurations are provided in Appendix C and D.
@@
Simulation Tasks ACT DP RDT Pi0 RDT +R1.0 Pi0 +R1.0 RDT +R2.0 Pi0 +R2.0
Stack Bowls Two 0.0% 0.0% 28.0% 41.0% 21.0% 29.0% 36.0% 55.0%
Pick Dual Bottles 0.0% 0.0% 5.0% 10.0% 13.0% 16.0% 20.0% 20.0%
Move Can Pot 0.0% 0.0% 20.0% 32.0% 23.0% 48.0% 25.0% 32.0%
Place Object Basket 0.0% 0.0% 11.0% 10.0% 14.0% 3.0% 17.0% 18.0%
Place Shoe 0.0% 0.0% 10.0% 12.0% 12.0% 7.0% 29.0% 24.0%
Average 0.0% 0.0% 14.8% 21.0% 16.6% 20.6% 25.4% 29.8%
As shown in Table 4, we observe that models fine-tuned with RoboTwin 1.0 data show negligible improvements in average success rate compared to their pretrained counterparts, indicating that data without domain randomization does not help the model handle environmental variations. In contrast, models pretrained with RoboTwin 2.0 data exhibit significantly improved generalization. Specifically, RDT and Pi0 achieve absolute gains of 10.6% and 8.8%, corresponding to relative improvements of 71.6% and 41.9%, respectively. Notably, this performance gain persists even though the downstream tasks were trained using only clean, non-randomized data. This suggests that domain-randomized pretraining with RoboTwin 2.0 effectively equips models with robustness to visual and spatial variations. As a result, models pretrained with RoboTwin 2.0 can adapt to new tasks without requiring additional data augmentation or complex scene variations.
@@ 4.4 Evaluation on Sim-to-Real Performance
To evaluate the effectiveness of RoboTwin 2.0 in enhancing real-world policy robustness, we conduct controlled experiments on four bimanual manipulation tasks: Stack Bowls, Handover Block, Pick Bottle, and Click Bell. The experiments use RDT as the policy backbone and are executed on the COBOT-Magic dual-arm robotic platform. For each task, we compare three training settings: (1) a baseline trained on 10 real-world demonstrations collected in clean tabletop environments; (2) the same real-world data augmented with 1,000 domain-randomized synthetic trajectories generated by RoboTwin 2.0 under cluttered scenes with diverse lighting and backgrounds; and (3) a synthetic-only setting trained solely on the 1,000 domain-randomized trajectories. To further enhance robustness to real-world jitter and calibration errors, we apply random 3D spatial perturbations to simulated object and robot poses, with the total displacement magnitude bounded by 1cm.
For evaluation, we define four test configurations: (i) clean tabletop with seen backgrounds (same as used in the 10 real-world demos); (ii) clean tabletop with unseen backgrounds; (iii) cluttered tabletop with seen backgrounds; and (iv) cluttered tabletop with unseen backgrounds, as illustrated in Fig.11. Notably, the synthetic-only (zero-shot) setting does not include seen backgrounds during training, so corresponding entries in the success rate table are left blank. This setup allows us to assess whether RoboTwin 2.0 can enable robust policy generalization without requiring additional real-world data from visually complex environments. Results are shown in Table 5.
Table 5: Real-World Experiment Results. We conduct controlled experiments on 4 dual-arm tasks: Bowls Stack, Block Handover, Bottle Pick, and Bell Click, each evaluated under 4 different settings.
@@
[…]
The experimental results show that real-world bimanual policies augmented with RoboTwin 2.0 data achieve substantial improvements across all evaluation settings. When adding 1,000 domain-randomized synthetic trajectories from RoboTwin 2.0, the success rates increased by 13.5%, 27.5%, 23.5%, and 33.0% across the four test configurations, respectively. For the zero-shot setting trained solely on synthetic data, we observe notable improvements of 21.0% and 20.5% on the two unseen-background scenarios. Notably, performance gains tend to increase with scene complexity, indicating that RoboTwin 2.0 enhances policy robustness under challenging visual conditions.
These gains can be attributed to two key factors. First, the visual and physical fidelity of RoboTwin 2.0 closely matches real-world conditions, enabling strong sim-to-real transfer. Second, even when real-world training data lacks such complexity, domain-randomized synthetic data effectively prepares the policy to handle challenging environments.
Interestingly, we also observe that models trained with RoboTwin 2.0 data exhibit consistent average success rates across difficulty levels in all four tasks. This suggests that the learned policies generalize well to different environments and are capable of selectively attending to task-relevant objects, further validating the effectiveness of training with RoboTwin 2.0.
@@
To evaluate the benchmarking utility and generalization challenges introduced by RoboTwin 2.0, we assess five policy models: ACT, DP, RDT, Pi0, and DP3. All evaluations are conducted on the full set of 50 benchmark tasks using the Aloha AgileX dual-arm embodiment. For each task, we collect 50 expert demonstrations in a clean simulation environment without domain randomization, which are used for training. During evaluation, each policy is tested with 100 rollouts under two conditions: a clean setting (Easy) and a domain-randomized setting (Hard). The Hard setting introduces cluttered scenes, random lighting, background textures, and varying tabletop heights. Detailed configuration settings are provided in Appendix. D. We report success rates as indicators of few-shot adaptation capability and robustness to challenging, randomized variations. Full benchmark results can be found at http://robotwin-platform.github.io/leaderboard/.
Table 6: RoboTwin 2.0 Simulation Benchmark (13 tasks sampled).
Simulation Task ACT DP DP3 RDT Pi0
@@
Grab Roller 66.0% 6.0% 98.0% 1.0% 77.0% 1.0% 74.0% 43.0% 96% 80.0%
Handover Mic 9.0% 0.0% 53.0% 0.0% 93.0% 7.0% 98.0% 41.0% 100% 13.0%
Lift Pot 7.0% 2.0% 37.0% 0.0% 85.0% 0.0% 82.0% 19.0% 84% 36.0%
Move Can Pot 0.0% 0.0% 42.0% 0.0% 28.0% 0.0% 47.0% 23.0% 74.0% 32.0%
Open Laptop 32.0% 0.0% 46.0% 0.0% 54.0% 3.0% 61.0% 36.0% 85.0% 46.0%
Pick Dual Bottles 4.0% 0.0% 26.0% 0.0% 25.0% 0.0% 42.0% 13.0% 57.0% 12.0%
Place Object Basket 0.0% 0.0% 17.0% 0.0% 2.0% 0.0% 42.0% 14.0% 62.0% 10.0%
Place Dual Shoes 0.0% 0.0% 7.0% 0.0% 1.0% 0.0% 4.0% 4.0% 15.0% 0.0%
Place Phone Stand 0.0% 0.0% 17.0% 0.0% 8.0% 1.0% 15.0% 6.0% 35.0% 7.0%
Put Bottles Dustbin 0.0% 0.0% 23.0% 0.0% 0.0% 0.0% 21.0% 4.0% 54.0% 13.0%
Put Object Cabinet 4.0% 18.0% 50.0% 17.0% 40.0% 16.0% 30.0% 30.0% 69.0% 29.0%
Stack Blocks Two 0.0% 0.0% 6.0% 0.0% 0.0% 0.0% 32.0% 1.0% 40.0% 1.0%
Stack Bowls Two 0.0% 0.0% 34.0% 0.0% 0.0% 0.0% 73.0% 21.0% 73.0% 41.0%
Average 9.4% 2.0% 35.1% 1.4% 31.8% 2.2% 47.8% 19.6% 64.9% 24.6%
Fig. 6 presents evaluation results on 13 representative tasks. As shown, non-pretrained models such as ACT, DP, and DP3 perform poorly under Hard conditions, with notably low success rates. In contrast, pretrained models like RDT and Pi0 show stronger resilience, suggesting that their vision-language-action pretraining provides useful priors for generalization to unseen scenarios. However, success rates still drop by 28.2% for RDT and 40.3% for Pi0 when transitioning from clean to randomized settings, indicating that robustness under heavy domain shifts remains a key challenge. This degradation likely stems from insufficient diversity in pretraining data, despite its large scale. Combined with the results from Figures 4.4 and 4.3, these findings highlight RoboTwin 2.0’s potential to complement existing pretraining datasets by providing diverse, domain-randomized trajectories that improve generalization and robustness under real-world variations.
@@ 5.1 Datasets and Benchmarks for Robotic Manipulation
Physics-based simulators form the backbone of modern manipulation research. SAPIEN xiang2020sapien supports fully dynamic interaction with over 2,300 articulated objects from PartNet-Mobility xiang2020sapien ; mo2019partnet , ManiSkill2 supplies more than four million demonstration frames across twenty task families maniskill2 , Meta-World metaworld offers fifty standardized manipulation tasks for multi-task and meta-reinforcement learning, CALVIN pairs long-horizon, language-conditioned instructions with rich sensor suites mees2022calvin , LIBERO liu2023libero defines 130 lifelong-learning tasks with high-quality human teleoperation data, and RoboVerse geng2025roboverse unifies multiple simulators and embodiments under a common benchmark with domain randomization.
In parallel, large-scale real-world datasets are essential to bridge simulation and reality. AgiBot World bu2025agibot provides over one million human-verified trajectories spanning 217 tasks in diverse deployment scenarios, RoboMIND wu2024robomind contributes 107 K teleoperated episodes (including annotated failures) across 479 tasks and four robot platforms, Open X-Embodiment openxembodiedment consolidates more than one million demonstrations from twenty-two robot embodiments into a unified format, and Bridge bridgedata delivers 60 K+ trajectories of 13 skills on low-cost hardware to evaluate transfer under varying visual and physical conditions.
RoboTwin-1.0 mu2025robotwin introduced a bidirectional twin framework that mirrors real teleoperated demonstrations with AI-generated simulated replicas for unified evaluation of dual-arm manipulation. In this work, we extend RoboTwin by integrating interactive LLM-driven feedback into the simulation loop and applying systematic domain randomization over visual, physical, and task parameters. These enhancements yield richer, more diverse training corpora and significantly improve policy robustness and generalization. A detailed comparison between RoboTwin 2.0 and prior datasets and benchmarks is provided in Appendix B.
@@ 5.2 Robot Learning in Manipulation
Many task-specific policy architectures wang2024rise ; ke20243d ; ze20243d ; chi2023diffusion ; fu2024mobile ; chen2025g3flow ; dexhanddiff ; wang2024rise ; adaptdiffuser ; liang2024skilldiffuser ; 10900471 ; wen2025dexvla achieve strong single-task performance but struggle to transfer across embodiments. In contrast, foundation models trained on million-scale, multi-robot corpora have enabled robust zero-shot generalization: RT-1 brohan2022rt-1 unifies vision, language and actions in a single transformer for real-time kitchen tasks; RT-2 brohan2023rt-2 co-fine-tunes large vision–language models on web and robot data to unlock semantic planning and object reasoning; diffusion-based RDT-1B liu2024rdt1b and the $\pi_{0}$black2024pi_0 capture diverse bimanual dynamics from over a million episodes. Vision–language–action (VLA) frameworks like OpenVLA openvla and CogACT li2024cogact , together with adaptations like Octo team2024octo , LAPA lapa , and OpenVLA-OFT openvla_oft demonstrate efficient fine-tuning to novel robots and sensor modalities.

### v2

@@ 4.2 Evaluating Efficiency with and without Adaptive Grasping
To evaluate the effectiveness of our embodiment-aware grasp augmentation strategy, we measure the task success rate of automated data collection across 50 RoboTwin 2.0 tasks on five different robot embodiments. As shown in Table 4.2, we compare our RoboTwin 2.0 pipeline against the RoboTwin 1.0 baseline, which lacks diverse grasping and candidate augmentation. Results show that our method improves success rates, particularly for robots with constrained planning spaces, achieving an average improvement of 8.3% across all embodiments. Specifically, for high-DoF arms with large reachable workspaces, such as Franka and UR5 (7-DoF), success rates remain largely unchanged, indicating limited benefit when the robot already has sufficient kinematic flexibility. However, for lower-DoF platforms such as Aloha-AgileX, Piper, and ARX-X5 (6-DoF), our method leads to substantial gains of 13.5%, 22.7%, and 5.6%, respectively. These results demonstrate that our approach provides additional feasible grasp options that effectively mitigate the planning limitations of low-DoF manipulators. Success rates for all tasks can be found in Appendix L.
@@
Our goal is to evaluate whether the domain-randomized data in RoboTwin 2.0 can endow models with robustness to environmental perturbations. To this end, we first pre-train RDT and Pi0 on 9,600 expert trajectories collected from 32 tasks (300 per task) under two settings: clean (non-randomized) and domain-randomized.
For comparison, we also evaluate the released pretrained weights of RDT and Pi0 without additional fine-tuning. To further study generalization, we select five unseen tasks and collect 50 clean demonstrations per task for single-task training and fine-tuning. Finally, all policies—including ACT, DP, RDT, and Pi0—are evaluated under domain-randomized conditions to measure robustness in previously unseen environments. Detailed configurations are provided in Appendix C and D.
@@
Simulation Tasks ACT DP RDT Pi0 RDT +Clean Pi0 +Clean RDT +Rand. Pi0 +Rand.
Stack Bowls Two 0.0% 0.0% 30.0% 41.0% 8.0% 55.0% 49.0% 62.0%
Pick Dual Bottles 0.0% 0.0% 13.0% 12.0% 12.0% 15.0% 17.0% 7.0%
Move Can Pot 4.0% 0.0% 12.0% 21.0% 13.0% 35.0% 18.0% 22.0%
Place Object Basket 0.0% 0.0% 17.0% 2.0% 9.0% 8.0% 6.0% 22.0%
Place Shoe 0.0% 0.0% 7.0% 6.0% 9.0% 6.0% 30.0% 18.0%
Open Laptop 0.0% 0.0% 32.0% 46.0% 21.0% 33.0% 35.0% 50.0%
Press Stapler 6.0% 0.0% 24.0% 29.0% 21.0% 26.0% 27.0% 31.0%
Turn Switch 2.0% 1.0% 15.0% 23.0% 24.0% 21.0% 16.0% 21.0%
Average 2.0% 0.0% 18.8% 22.5% 14.6% 24.9% 24.8% 29.1%
As shown in Table 3, we observe that models fine-tuned with clean data show negligible improvements in average success rate compared to their pretrained counterparts, indicating that data without domain randomization does not help the model handle environmental variations. This also suggests that the low success rate of pretrained VLAs in simulation is not due to a Real-to-Sim gap, since we provide clean simulation data yet observe no clear improvement. In contrast, models pretrained with RoboTwin 2.0 data exhibit significantly improved generalization. Specifically, RDT and Pi0 achieve relative improvements of 31.9% and 29.3%. Notably, this performance gain persists even though the downstream tasks were trained using only clean, non-randomized data. This demonstrates that domain-randomized pretraining with RoboTwin 2.0 effectively equips models with robustness to visual and spatial variations. As a result, models pretrained with RoboTwin 2.0 can adapt to new tasks without requiring additional data augmentation or complex scene variations.
@@ 4.4 Evaluation on Sim-to-Real Performance
To evaluate RoboTwin 2.0’s effectiveness in enhancing real-world policy robustness, we conduct experiments on four bimanual tasks: Stack Bowls, Handover Block, Pick Bottle, and Click Bell. All experiments use RDT as the policy backbone and are executed on the COBOT-Magic dual-arm platform. We compare three training settings: (1) 10 real-world demonstrations in clean tabletop environments; (2) the same demonstrations augmented with 1,000 domain-randomized synthetic trajectories generated under cluttered scenes with varied lighting and backgrounds; and (3) a synthetic-only setting trained solely on the 1,000 domain-randomized trajectories. To further improve robustness to camera jitter and calibration errors, we apply random 3D perturbations to simulated camera poses (position and orientation), with the displacement magnitude bounded by 1 cm.
Evaluation is conducted under four test configurations: (i) clean tabletop with seen backgrounds, (ii) clean tabletop with unseen backgrounds, (iii) cluttered tabletop with seen backgrounds, and (iv) cluttered tabletop with unseen backgrounds (Fig.10). Since the synthetic-only setting excludes seen backgrounds during training, the corresponding entries in Table 4 are omitted. This setup directly tests whether RoboTwin 2.0 enables robust policy generalization without additional real-world data from visually complex environments.
Table 4: Real-World Experiment Results. We conduct controlled experiments on 4 dual-arm tasks: Stack Bowls, Handover Block, Pick Bottle, and Click Bell, each evaluated under 4 different settings.
@@
Average Seen False 29.5% $\textbf{43.0\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+13.5\%}}$ /
True 14.0% $\textbf{41.5\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+27.5\%}}$ /
Unseen False 15.5% $\textbf{39.0\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+23.5\%}}$ ${36.5\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+21.0\%}}$
True 9.0% $\textbf{42.0\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+33.0\%}}$ ${29.5\%}_{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+20.5\%}}$
The experimental results show that real-world bimanual policies augmented with RoboTwin 2.0 achieve clear gains in robustness. In the few-shot setting—where 1,000 domain-randomized synthetic trajectories are combined with just 10 real-world demonstrations—the average success rate across all evaluation settings improves by 24.4%, with per-configuration gains of 13.5%, 27.5%, 23.5%, and 33.0%, respectively. In the zero-shot setting trained solely on synthetic data, we still observe notable improvements of 21.0% and 20.5% on the two unseen-background scenarios. Notably, performance gains become larger in visually complex scenes, indicating that RoboTwin 2.0 is especially effective under challenging conditions.
These improvements stem from two factors: (1) the high visual and physical fidelity of RoboTwin 2.0, enabling direct sim-to-real transfer, and (2) the ability of domain-randomized synthetic data to prepare policies for environmental variations absent from clean real-world demonstrations. Importantly, the strong performance of the few-shot setting suggests that only minimal real-world data is needed to effectively bridge the sim-to-real gap.
@@
To evaluate the benchmarking utility and generalization challenges of RoboTwin 2.0, we assess five policy models: ACT, DP, RDT, Pi0, and DP3. All VLAs are fine-tuned from their released pretrained weights in the single-task setting. Evaluations are conducted on all 50 benchmark tasks using the Aloha AgileX dual-arm embodiment. For each task, 50 clean expert demonstrations are used for training, and policies are tested with 100 rollouts under two conditions: Easy (clean) and Hard (domain-randomized with clutter, lighting, textures, and height variations). We provide a visualization of the benchmark setting in Appendix J. We report success rates as indicators of few-shot adaptation and robustness. Detailed setups are provided in Appendix C and D, and full results are available in Appendix K and on the Leaderboard.
Table 5: Subset of RoboTwin 2.0 benchmark. Full results in Appendix K and Leaderboard.
Simulation Task RDT Pi0 ACT DP DP3
@@
Adjust Bottle 81% 75% 90% 56% 97% 23% 97% 0% 99% 3%
Beat Block Hammer 77% 37% 43% 21% 56% 3% 42% 0% 72% 8%
Blocks Ranking RGB 3% 0% 19% 5% 1% 0% 0% 0% 3% 0%
Blocks Ranking Size 0% 0% 7% 1% 0% 0% 1% 0% 2% 0%
Click Alarmclock 61% 12% 63% 11% 32% 4% 61% 5% 77% 14%
Click Bell 80% 9% 44% 3% 58% 3% 54% 0% 90% 0%
Dump Bin Bigbin 64% 32% 83% 24% 68% 1% 49% 0% 85% 53%
Grab Roller 74% 43% 96% 80% 94% 25% 98% 0% 98% 2%
Handover Block 45% 14% 45% 8% 42% 0% 10% 0% 70% 0%
Handover Mic 90% 31% 98% 13% 85% 0% 53% 0% 100% 3%
Hanging Mug 23% 16% 11% 3% 7% 0% 8% 0% 17% 1%
Lift Pot 72% 9% 84% 36% 88% 0% 39% 0% 97% 0%
Move Pillbottle Pad 8% 0% 21% 1% 0% 0% 1% 0% 41% 0%
Average (in %) 34.5 13.7 46.4 16.3 29.7 1.7 28.0 0.6 55.2 5.0
Fig.5 and Appendix K report results on 50 tasks. Non-pretrained models (ACT, DP, DP3) perform poorly under Hard conditions, while pretrained models (RDT, Pi0) show stronger resilience, suggesting that vision–language–action pretraining provides useful priors for generalization. Still, success rates drop by 20.8% (RDT) and 30.1% (Pi0) from clean to randomized settings, underscoring robustness under domain shifts as a key challenge, likely due to limited diversity in pretraining data. DP3 achieves the best few-shot performance with limited samples, highlighting the role of 3D information, though its strong results partly stem from perfect point clouds and clean background segmentation in simulation. Together with Sections 4.3 and 4.4, these findings show RoboTwin 2.0’s value in complementing existing datasets with diverse, domain-randomized trajectories for improved generalization and robustness.
@@ 5.1 Datasets and Benchmarks for Robotic Manipulation
Physics-based simulators underpin modern manipulation research. Existing platforms provide complementary strengths: SAPIEN xiang2020sapien enables dynamic interaction with 2,300+ articulated objects; ManiSkill2 maniskill2 supplies millions of demonstrations; Meta-World metaworld , CALVIN mees2022calvin , LIBERO liu2023libero , and RoboVerse geng2025roboverse target multi-task, language-conditioned, lifelong, and domain-randomized settings; RoboCasa robocasa2024 offers large-scale human demonstrations but lacks automation and dual-arm focus.
Large-scale real-world datasets further bridge sim-to-real: AgiBot World bu2025agibot , RoboMIND wu2024robomind , Open X-Embodiment openxembodiedment , and Bridge bridgedata contribute millions of trajectories across diverse tasks, robots, and environments.
RoboTwin-1.0 mu2025robotwin mirrored real demonstrations with simulated replicas for dual-arm benchmarking. In this work, RoboTwin 2.0 integrates LLM-driven feedback and systematic domain randomization across visual, physical, and task dimensions, producing richer corpora that improve policy robustness and generalization. A detailed comparison with prior benchmarks is provided in Appendix B.
@@ 5.2 Robot Learning in Manipulation
Many task-specific policy architectures wang2024rise ; ke20243d ; ze20243d ; chi2023diffusion ; fu2024mobile ; chen2025g3flow ; dexhanddiff ; wang2024rise ; adaptdiffuser ; liang2024skilldiffuser ; 10900471 ; wen2025dexvla ; chen2025benchmarking achieve strong single-task performance but struggle to transfer across embodiments. In contrast, foundation models trained on million-scale, multi-robot corpora have enabled robust zero-shot generalization: RT-1 brohan2022rt-1 unifies vision, language and actions in a single transformer for real-time kitchen tasks; RT-2 brohan2023rt-2 co-fine-tunes large vision–language models on web and robot data to unlock semantic planning and object reasoning; diffusion-based RDT-1B liu2024rdt1b and the $\pi_{0}$black2024pi_0 capture diverse bimanual dynamics from over a million episodes. Vision–language–action (VLA) frameworks like OpenVLA openvla and CogACT li2024cogact , together with adaptations like Octo team2024octo , LAPA lapa , and OpenVLA-OFT openvla_oft demonstrate efficient fine-tuning to novel robots and sensor modalities.